There are a number of matrix options in AMS-TEX. There is pmatrix.

A = ( \begin{matrix}x − λ & 1 & 0 \\ 0 & x − λ & 1 \\ 0 & 0 & x − λ \end{matrix} )

There is bmatrix.

A = [ \begin{matrix}x − λ & 1 & 0 \\ 0 & x − λ & 1 \\ 0 & 0 & x − λ \end{matrix} ]

There is vmatrix.

A = | \begin{matrix}x − λ & 1 & 0 \\ 0 & x − λ & 1 \\ 0 & 0 & x − λ \end{matrix} |

There is Vmatrix.

A = ∥ \begin{matrix}x − λ & 1 & 0 \\ 0 & x − λ & 1 \\ 0 & 0 & x − λ \end{matrix} ∥

Justification and other formatting is available.

A = ( \begin{matrix}cos θ & sin θ \\ − sin θ & cos θ \end{matrix} )

And, AMS-TEX supplies it own cases command.

f(x) = \{ \begin{matrix}x + 1, & if x > 0 \\ x − 1, & if x ≤ 0 \end{matrix}